A 56 kg skier starts from rest at the top of a 1200-m-long trail which drops a total of 160 m from top to bottom. At the bottom, the skier is moving 11.0 m/s .

How much energy was dissipated by friction?

Express your answer using two significant figures.
Answer #1

Total mechanical energy of skier = mgh = (56)(9.8)(160) = 87,808 J

KE of skier at bottom of trail = 1/2mV² = (0.5)(56)(11)² = 3388 J

Energy dissipated by friction = 87,808 - 3388 = 84,420 ≈ 84,000 J ANS